OpenStudy (hnf1109):

Given the relation y= -4x squared - 8, if the input is 2, what is the output? A. 8 B. -8 C. -16 D. -24

OpenStudy (jango_in_dtown):

So we have y= -4x^2-8 for x=2, y=-4(2)^2-8=-16-8=-24

OpenStudy (hnf1109):

Thankyou!!